OpenStudy (surryyy):

17! / 4! x 13! Answers: .0238 23,800 2,380 Some number less than 1

OpenStudy (anonymous):

17! = 17 x16 x15x...x1 but there is a button on your calculator that will do this also, so ust type 1in 17 and hit the factoria button ( which is normally an n followed by an exclamation mark)

OpenStudy (anonymous):

Alternativey, you coud note that 17! = 17 x 16 x15 x 14 x 13 x12...etc but 13! aso has 13x 12 ...etc so these would cancel each other to leave: 17 x 16 x 15 x 14/3!

OpenStudy (surryyy):

This is so confusing..

OpenStudy (anonymous):

what is?

OpenStudy (surryyy):

this whole thing.

OpenStudy (anonymous):

these are known as factorials, in simpe terms it means if I have 17! (which you pronounce as "17 factorial") it means multiply 17 x 16 x 15 etc all the wway down to 1. Nothing more to it. Its a shorthand way of saying this.

OpenStudy (surryyy):

ok but i cant find my final answer.

OpenStudy (anonymous):

17 x16 x15 x14 = 57120 Now to divide by 4! = 4 x 3 x 2 x 1 = 24 so we get 57120 /24 = 2380

OpenStudy (surryyy):

oh ok. thnx.

OpenStudy (uri):

@surryyy son

OpenStudy (surryyy):

im a girl

OpenStudy (uri):

Idc,i am here to help you.

OpenStudy (surryyy):

lol ok. thnx

OpenStudy (uri):

It's just that 17! =17*16*15*14*13*12*11*10*9*8*7*6*5*4*3*2*1 ============================= = 4!*13! 4*3*2*1*13*12*11*10**9*8*7*6*5*4*3*2*1

OpenStudy (uri):

Now,cancel from 13 to 1 from both numerator and denominator.

OpenStudy (uri):

17*16*15*14 ========= = 4*3*2*1

OpenStudy (uri):